OpenStudy (anonymous):

Question asks to convert the polar equation to a rectangular equation and express in standard form

OpenStudy (anonymous):

r=3sec(theta)

OpenStudy (tkhunny):

Oh, you can't let these easy, fun ones get away! \(r = 3\cdot\sec(\theta) = \dfrac{3}{\cos(\theta)}\) Do you see it, yet?

OpenStudy (anonymous):

yea the problem im having is expressing in standard form, i alredy have that x=3 because rcostheta becomes x

OpenStudy (tkhunny):

That's it. Maybe x - 3 = 0? Why does anyone care what form it's in? So, what about my premise? Was it fun when you saw it?

OpenStudy (anonymous):

lol my professor cares although he really dint go over how to change it into standard form

OpenStudy (tkhunny):

I always forget Standard and Normal. They are relatively useless for now. The better forms are the ones that give you information. ax + by = c ax + by + c = 0 Make sure you know the defintions. It's a little less obvious that it is right when a or b is zero, such as your case.

OpenStudy (anonymous):

thank you